# On $f$-Divergence Principled Domain Adaptation: An Improved Framework

### v1

@@ 3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
Acuna et al. (2021) also applies Rademacher complexity-based bound to further upper bound $\lambda^{*}$ by its empirical version, namely $\hat{\lambda}^{*}=\min_{h^{*}\in\mathcal{H}}R_{\hat{\mu}}(h^{*})+R_{\hat{\nu}}(h^{*})$. However, since there is no target label available, $R_{\hat{\nu}}(h^{*})$ is still uncomputable, invoking $\hat{\lambda}^{*}$ here has no clear advantage.
@@ 4 New $f$-Divergence-based DA Theory
While $\widetilde{\rm D}_{\phi}^{h,\mathcal{H}}$ serves as a valid domain discrepancy measure in DA theory, it exaggerates the domain difference without appropriate control. It’s noteworthy that Acuna et al. (2021) attempts to demonstrate their absolute discrepancy is upper bounded by ${\rm D}_{\phi}$ (Acuna et al., 2021, Lemma 1), but this is problematic; as $\sup{U}\geq 0$ does not imply $\sup{U}=\sup\left|U\right|$ when $U$ is not a positive function. When designing their $f$-DAL algorithm, they drop the absolute value function in their hypothesis-specified $f$-divergence (see Eq. (6)). Consequently, the remarkable performance of $f$-DAL reveals a significant gap from their theoretical foundation.
@@
If $\mathcal{H}^{\ell}=\mathcal{G}$, ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ is an affine transformation of Lemma 2.1 (i.e. $g\to tg+\alpha$) and a scaling transformation of Lemma 2.2. Importantly, unlike $\widetilde{\rm D}_{\phi}^{h,\mathcal{H}}$, it’s easy to see that ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\leq{\rm D}_{\phi}(\nu||\mu)$.
Our ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ retains some common properties of the discrepancies defined in the DA theory literature. First, as $t=0$ leads to ${{\mathbb{E}_{\nu}}\left[t\cdot\ell(h,h^{\prime})\right]}=I_{\phi,\mu}^{h}(t\ell\circ h^{\prime})=0$, the non-negativity of ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ is immediately justified. In addition, its asymmetric property is also preferred in DA, as discussed in the previous works (Zhang et al., 2019, 2020). Moreover, when $\mu=\nu$, by the definition of $\phi^{*}$, we have ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)=0$.
@@
Let $\psi(x)\triangleq\phi(x+1)$, and $\psi^{*}$ is its convex conjugate. Define $K_{h^{\prime},\mu}(t)\triangleq\inf_{\alpha}{{\mathbb{E}_{\mu}}\left[\psi^{*}(t\cdot\ell(h,h^{\prime})+\alpha)\right]}$. Let $K_{\mu}(t)=\sup_{h^{\prime}\in\mathcal{H}}K_{h^{\prime},\mu}(t)$, then for any $h,h^{\prime}\in\mathcal{H}$,
@@
It is worth reminding that both $K_{\mu}$ and $K^{*}_{\mu}$ still depend on $h$, although we ignore $h$ in the notations to avoid cluttering. We are now ready to give a target error bound.
@@
$*.\displaystyle R_{\nu}(h)\leq R_{\mu}(h)+\inf_{t\geq 0}\frac{\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)+K_{\mu}(t)}{t}+\lambda^{*}.$ (3)
Furthermore, let $\ell\in[0,1]$, if $\phi$ is twice differentiable and ϕ′′superscriptitalic-$\phi^{\prime\prime}$ is monotone, then
$*.\displaystyle R_{\nu}(h)\leq R_{\mu}(h)+\sqrt{\frac{2}{\phi^{\prime\prime}(1)}{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}+\lambda^{*}.$ (4)
As an application of Eq. (4), we consider the case of KL, namely ${\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)$. We have the following result.
@@
Let $\ell\in[0,1]$, then for any $h\in\mathcal{H}$, we have
$*.\displaystyle R_{\nu}(h)\leq R_{\mu}(h)+\sqrt{2{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)}+\lambda^{*}.$ (5)
As ${\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)\leq{\mathrm{D}}_{\rm KL}(\nu||\mu)$, the bound in Wang & Mao (2023a, Theorem 4.2) can be recovered by Corollary 4.1 for the same bounded loss function. We also remark that the boundedness assumption in Corollary 4.1 can be further relaxed by applying the same sub-Gaussian assumption as in Wang & Mao (2023a, Theorem 4.2).
@@
${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\leq{\rm D}^{h,\mathcal{H}}_{\phi}$ $(\hat{\nu}||\hat{\mu})\!+\!2\left|t_{0}\right|\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})\!+\!2L\left|t_{0}\right|\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})$
$+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}+\sqrt{\frac{\log(1/\delta)}{m}}\right).$
@@
The function $\phi^{*}$ needs not to be globally Lipschitz continues; it can be locally Lipschitz for a bounded domain. For example, in the case of KL, $\phi^{*}$ is $e$-Lipschitz continues when $\ell\in[0,1]$. Moreover, although $t_{0}$ might not always have a closed-form expression, in the case of $\chi^{2}$-DD, we know that $t_{0}=\frac{2\left({{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime*})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime*})\right]}\right)}{\mathrm{Var}_{\mu}\left(\ell(h,h^{\prime*})\right)}$ where $h^{\prime*}$ is the corresponding optimal hypothesis.
It is also straightforward to obtain concentration results for $K_{\mu}(t)-K_{\hat{\mu}}(t)$ and $R_{\mu}(h)-R_{\hat{\mu}}(h)$. Substituting these results into Eq. (3) yields the final generalization bound based on $f$-DD. Or alternatively, one can directly substitute Lemma 4.2 into Eq. (4) (See Appendix D.6). However, in this case, the empirical $f$-DD and other terms in Lemma 4.2 will feature a square root, slowing down the convergence rate. We address this limitation in the following section.
@@ 5 Sharper Bounds via Localization
A localization technique in DA theory is recently studied in Zhang et al. (2020). We now incorporate it into our framework. First, we define a localized hypothesis space, formally referred to as the (true) Rashomon set (Breiman, 2001; Semenova et al., 2022; Fisher et al., 2019).
@@
Notice that the Rashomon set $\mathcal{H}_{r}$ implicitly depends on the data distribution. In this paper, we specifically define $\mathcal{H}_{r}$ by the source domain distribution $\mu$. Then, we define our localized $f$-DD below.
@@
As also mentioned at the end of the previous section, Eq. (4) of Theorem 4.1 (and Eq. (5) of Corollary 4.1), involves a square root function for $f$-DD, potentially indicative of a slow-rate bound (e.g., if $\mathrm{D}^{h,\mathcal{H}}_{\phi}\in\mathcal{O}(1/n)$, then the bound decays with $\mathcal{O}(1/\sqrt{n})$). We now show that how the localized $f$-DD achieves a fast-rate error bound.
@@
By the triangle property, $R^{r}_{\mu}(h)\leq r+r_{1}$. In this case, a small $r_{1}$ will reduce both the first term and the third term in the bound. However, determining the optimal value for $r$ is intricate. On the one hand, we hope $r$ is small so that ${\mathrm{D}}^{h,\mathcal{H}_{r}}_{\phi}(\nu||\mu)$ and $R^{r}_{\mu}(h)$ are both small. On the other hand, if $r$ is too small, specifically if $r<\lambda^{*}$, then it’s possible that $\lambda_{r}^{*}>\lambda^{*}$ because the optimal hypothesis minimizing the joint risk may not exist in $\mathcal{H}_{r}$. Additionally, if both $r_{1}$ and $r$ are too small, the effective space for optimizing $C_{1}$ and $C_{2}$ may also be limited. Therefore, the value of $r$ engenders a complex trade-off among the three terms.
@@
Combining Theorem 5.1 with Lemma 4.2 and following routine steps will obtain the generalization bound based on $f$-DD, where the local Rademacher complexity (Bartlett et al., 2005) will be involved. However, one may feel unsatisfied without an explicit clue for the condition $K_{h^{\prime},\mu}(C_{1})\leq C_{1}C_{2}{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$ in Theorem  5.1. In fact, exploring concentration results under this condition is a central theme in obtaining fast-rate PAC-Bayesian generalization bounds (Catoni, 2007; Seldin et al., 2012; Tolstikhin & Seldin, 2013; Yang et al., 2019; Alquier, 2021) and the information-theoretic generalization bounds (Hellström & Durisi, 2021, 2022; Wang & Mao, 2023c, b). Building upon similar ideas from these works, we now establish a sharper generalization result for our localized KL-DD measure, where the fast-rate condition is more explicit.
One key ingredient is the following result.
@@
Let $\ell\in[0,1]$, and let $C_{1}>0$ and $C_{2}\in(0,1)$ satisfy $\left(e^{C_{1}}-C_{1}-1\right)\left(1+C_{2}^{2}\min\{r_{1}+r,1\}\right)-tC_{2}\leq 0$. Then, for any $h\in\mathcal{H}_{r_{1}}$ and $h^{\prime}\in\mathcal{H}_{r}$, we have
@@
As an extreme case, if $r+r_{1}\to 0$, then we let $C_{2}\to 1$, the condition in the lemma indicates that $C_{1}<1.26$. Hence, the optimal bound becomes ${{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime})\right]}\leq 0.79\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}({\nu}||{\mu})$. This bound remains valid even without $r+r_{1}\to 0$. It holds when the Rashomon set $H_{r}$ is “consistent” with a given $h$, meaning all hypotheses in $H_{r}$ have similar predictions to $h$ on the source domain data. As an another case, if $r+r_{1}\geq 1$ and $\sup_{h^{\prime}}{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$ is also large, we may prefer a small $C_{2}$, such as setting $C_{2}=0.1$. This suggests $C_{1}<0.2$, resulting in the optimal bound ${{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}\leq 5\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}({\nu}||{\mu})+0.1{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$. The conditions in Lemma 5.1 are common in many fast-rate bound literature, such as Hellström & Durisi (2022, Theorem 3).
@@
$R_{\nu}$ $(h)\leq R_{\hat{\mu}}(h)\!+\!\frac{\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}(\hat{\nu}||\hat{\mu})}{C_{1}}\!+\!C_{2}R^{r}_{\mu}(h)\!+\!\mathcal{O}\left(\hat{\mathfrak{R}}_{\mathcal{T}}\left(\mathcal{H}^{\ell}_{r}\right)\right.$
+^$+\left.\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{\max\{r,r_{1}\}}\right)\right)\!+\!\mathcal{O}\left(\frac{\log(1/\delta)}{n}\!+\!\frac{\log(1/\delta)}{m}\right)$
$+\mathcal{O}\left(\sqrt{\frac{(r_{1}+r)\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{m}}\right)+\lambda_{r}^{*}.$
@@
Due to the non-negativity of $f$-DD, a similar generalization bound also applies to the Jeffereys divergence (or symmetrized KL divergence) (Jeffreys, 1946) counterpart, which is simply the sum of KL divergence and reverse KL divergence (i.e. $\mathrm{D}_{\rm KL}(\hat{\mu}||\hat{\nu})+\mathrm{D}_{\rm KL}(\hat{\nu}||\hat{\mu})$). Furthermore, considering the fact that $\mathrm{D}_{\rm KL}(\nu||\mu)\leq\log(1+\chi^{2}(\nu||\mu))\leq\chi^{2}(\nu||\mu)$ (Polyanskiy & Wu, 2023), one might anticipate a similar bound for $\chi^{2}$-DD, which we defer to Appendix D.10.
This generalization bound suggests that when $r+r_{1}$ is small, not only are the first four terms (including the local Rademacher complexities) reduced, but it also causes $\mathcal{O}\left(\frac{1}{n}+\frac{1}{m}\right)$ to dominate the convergence rate of the bound. In a practical algorithm, when empirical risk of source domain is always minimized to zero (i.e. the realizable case), then $r_{1}$ itself may have a fast vanishing rate (e.g., $\mathcal{O}(1/n)$).

### v2

@@ 3 Warm-Up: Refined Absolute $f$-Divergence Domain Discrepancy
[1] also applies Rademacher complexity-based bound to further upper bound $\lambda^{*}$ by its empirical version, namely $\hat{\lambda}^{*}=\min_{h^{*}\in\mathcal{H}}R_{\hat{\mu}}(h^{*})+R_{\hat{\nu}}(h^{*})$. However, since there is no target label available, $R_{\hat{\nu}}(h^{*})$ is still not computable, invoking $\hat{\lambda}^{*}$ here has no clear advantage.
@@ 4 New $f$-Divergence-Based DA Theory
While $\widetilde{\rm D}_{\phi}^{h,\mathcal{H}}$ serves as a valid domain discrepancy measure in DA theory, it exaggerates the domain difference without appropriate control. It’s noteworthy that [1] attempts to demonstrate their discrepancy with the absolute value function is upper bounded by ${\rm D}_{\phi}$ [1, Lemma 1], but this is problematic; as $\sup{U}\geq 0$ does not imply $\sup{U}=\sup\left|U\right|$ when $U$ is not a positive function. Note that the error in [1, Lemma 1] is also identified in [35]. Furthermore, when designing their $f$-DAL algorithm, they drop the absolute value function in their hypothesis-specified $f$-divergence (see Eq. (5)). Consequently, the remarkable performance of $f$-DAL reveals a significant gap from their theoretical foundation.
@@
If $\mathcal{H}^{\ell}=\mathcal{G}$, then ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ is an affine transformation of Lemma 2.1 (i.e. $g\to tg+\alpha$) and a scaling transformation of Lemma 2.2. Importantly, unlike $\widetilde{\rm D}_{\phi}^{h,\mathcal{H}}$, it’s easy to see that ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\leq{\rm D}_{\phi}(\nu||\mu)$.
Our ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ retains some common properties of the discrepancies defined in the DA theory literature. First, as $t=0$ leads to ${{\mathbb{E}_{\nu}}\left[t\cdot\ell(h,h^{\prime})\right]}=I_{\phi,\mu}^{h}(t\ell\circ h^{\prime})=0$, the non-negativity of ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)$ is immediately justified. In addition, its asymmetric property is also preferred in DA, as discussed in the previous works [9, 10]. Moreover, when $\mu=\nu$, by the definition of $\phi^{*}$, we have ${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)=0$.
@@
Let $\psi(x)\triangleq\phi(x+1)$, and $\psi^{*}$ is its convex conjugate. For any $h^{\prime},h\in\mathcal{H}$ and $t\in\mathbb{R}$, define $K_{h^{\prime},\mu}(t)\triangleq\inf_{\alpha}{{\mathbb{E}_{\mu}}\left[\psi^{*}(t\cdot\ell(h,h^{\prime})+\alpha)\right]}$. Let $K_{\mu}(t)=\sup_{h^{\prime}\in\mathcal{H}}K_{h^{\prime},\mu}(t)$, then for any $h,h^{\prime}\in\mathcal{H}$,
@@
It is worth reminding that $K_{\mu}$ and $K^{*}_{\mu}$ both depend on $h$, although we ignore $h$ in the notations to avoid cluttering. We are now ready to give a target error bound based on our $f$-DD.
@@
$R_{\nu}(h)\leq R_{\mu}(h)+\inf_{t\geq 0}\frac{\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)+K_{\mu}(t)}{t}+\lambda^{*}.$ (3)
Furthermore, let $\ell\in[0,1]$, if $\phi$ is twice differentiable and $\phi^{\prime\prime}$ is monotone, then
$R_{\nu}(h)\leq R_{\mu}(h)+\sqrt{\frac{2}{\phi^{\prime\prime}(1)}{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}+\lambda^{*}.$ (4)
The following is an application of Eq. (4), where we consider the case of KL, namely ${\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)$.
@@
Let $\ell\in[0,1]$, then for any $h\in\mathcal{H}$, we have $R_{\nu}(h)\leq R_{\mu}(h)+\sqrt{2{\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)}+\lambda^{*}$.
As ${\mathrm{D}}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)\leq{\mathrm{D}}_{\rm KL}(\nu||\mu)$, the bound in [19, Theorem 4.2] can be recovered by Corollary 4.1 for the same bounded loss function. We also remark that the boundedness assumption in Corollary 4.1 can be further relaxed by applying the same sub-Gaussian assumption as in [19, Theorem 4.2].
@@
${\rm D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\leq{\rm D}^{h,\mathcal{H}}_{\phi}(\hat{\nu}||\hat{\mu})+2\left|t_{0}\right|\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})+2L\left|t_{0}\right|\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}+\sqrt{\frac{\log(1/\delta)}{m}}\right).$
@@
The function $\phi^{*}$ needs not to be globally Lipschitz continuous; it can be locally Lipschitz for a bounded domain. For example, in the case of KL, $\phi^{*}$ is $e$-Lipschitz continuous when $\ell\in[0,1]$. Moreover, although the distribution-dependent quantity $t_{0}$ might not always have a closed-form expression, in the case of $\chi^{2}$-DD, we know that $t_{0}=\frac{2\left({{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime*})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime*})\right]}\right)}{\mathrm{Var}_{\mu}\left(\ell(h,h^{\prime*})\right)}$, where $h^{\prime*}$ is the corresponding optimal hypothesis.
It is also straightforward to obtain concentration results for $K_{\mu}(t)-K_{\hat{\mu}}(t)$ and $R_{\mu}(h)-R_{\hat{\mu}}(h)$. Substituting these results into Eq. (3) obtains the final generalization bound based on $f$-DD. Or alternatively, one can directly substitute Lemma 4.2 into Eq. (4) (See Appendix D.7). However, in this case, the empirical $f$-DD and other terms in Lemma 4.2 will feature a square root, slowing down the convergence rate. We address this limitation in the following section.
@@ 5 Sharper Bounds via Localization
A localization technique in DA theory is recently studied in [10]. We now incorporate it into our framework with some novel applications. First, we define a localized hypothesis space, formally referred to as the (true) Rashomon set [36, 37, 38].
@@
Notice that the Rashomon set $\mathcal{H}_{r}$ implicitly depends on the data distribution. In this paper, we specifically define $\mathcal{H}_{r}$ by the source domain distribution $\mu$. Then, we define our localized $f$-DD:
@@
As also mentioned at the end of the previous section, Eq. (4) of Theorem 4.1 (and Corollary 4.1), involves a square root function for $f$-DD, potentially indicative of a slow-rate bound (e.g., if $\mathrm{D}^{h,\mathcal{H}}_{\phi}\in\mathcal{O}(1/n)$, then the bound decays with $\mathcal{O}(1/\sqrt{n})$). We now show that how the localized $f$-DD achieves a fast-rate error bound.
@@
By the triangle property, $R^{r}_{\mu}(h)\leq r+r_{1}$. In this case, a small $r_{1}$ will reduce both the first term and the third term in the bound. However, determining the optimal value for $r$ is intricate. On the one hand, we hope $r$ is small so that ${\mathrm{D}}^{h,\mathcal{H}_{r}}_{\phi}(\nu||\mu)$ and $R^{r}_{\mu}(h)$ are both small. On the other hand, if $r$ is too small, specifically if $r<\lambda^{*}$, then it’s possible that $\lambda_{r}^{*}>\lambda^{*}$ because the optimal hypothesis minimizing the joint risk may not exist in $\mathcal{H}_{r}$. Additionally, if both $r_{1}$ and $r$ are too small, the effective space for optimizing $C_{1}$ and $C_{2}$ may also be limited. Therefore, the value of $r$ involves a complex trade-off among the three terms.
@@
Combining Theorem 5.1 with Lemma 4.2 and following routine steps will obtain the generalization bound based on $f$-DD, where the local Rademacher complexity [24] will be involved. However, one may feel unsatisfied without an explicit clue for the condition $K_{h^{\prime},\mu}(C_{1})\leq C_{1}C_{2}{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$ in Theorem  5.1. In fact, exploring concentration results under this condition is a central theme in obtaining fast-rate PAC-Bayesian generalization bounds [25, 39, 40, 41, 26] and the information-theoretic generalization bounds [27, 28, 29, 42]. Building upon similar ideas from these works, we now establish a sharper generalization result for our localized KL-DD measure, where the fast-rate condition is more explicit. One key ingredient is the following result.
@@
Let $\ell\in[0,1]$, and let the constants $C_{1}>0$ and $C_{2}\in(0,1)$ satisfy the condition $\left(e^{C_{1}}-C_{1}-1\right)\left(1-\min\{r_{1}+r,1\}+C_{2}^{2}\min\{r_{1}+r,1\}\right)\leq C_{1}C_{2}$. Then, for any $h\in\mathcal{H}_{r_{1}}$ and $h^{\prime}\in\mathcal{H}_{r}$, we have
@@
As an extreme case, if $r+r_{1}\to 0$ (implying ${{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}=0$), then let $C_{2}\to 1$, the condition in the lemma indicates that $C_{1}<1.26$. Hence, the optimal bound becomes ${{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime})\right]}\leq 0.79\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}({\nu}||{\mu})$. This bound remains valid even without $r+r_{1}\to 0$. It holds when the Rashomon set $H_{r}$ is “consistent” with a given $h$, meaning all hypotheses in $H_{r}$ have similar predictions to $h$ on the source domain data. As an another case, if $r+r_{1}\geq 1$ and $\sup_{h^{\prime}}{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$ is also large, we may prefer a small $C_{2}$, such as setting $C_{2}=0.1$. In this case, the condition becomes $\left(e^{C_{1}}-C_{1}-1\right)C_{2}\leq C_{1}$, suggesting that $C_{1}<3.74$. This results in the optimal bound ${{\mathbb{E}_{\nu}}\left[\ell(h,h^{\prime})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}\leq 0.27\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}({\nu}||{\mu})+0.1{{\mathbb{E}_{\mu}}\left[\ell(h,h^{\prime})\right]}$. The condition in Lemma 5.1 is common in many fast-rate bound literature, such as [28, Theorem 3].
@@
$R_{\nu}(h)\leq$ $R_{\hat{\mu}}(h)+\frac{\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}(\hat{\nu}||\hat{\mu})}{C_{1}}+C_{2}R^{r}_{\mu}(h)+\mathcal{O}\left(\hat{\mathfrak{R}}_{\mathcal{T}}\left(\mathcal{H}^{\ell}_{r}\right)+\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{\max\{r,r_{1}\}}\right)\right)$
$+\mathcal{O}\left(\frac{\log(1/\delta)}{n}+\frac{\log(1/\delta)}{m}\right)+\mathcal{O}\left(\sqrt{\frac{(r_{1}+r)\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{m}}\right)+\lambda_{r}^{*}.$
@@
Due to the non-negativity of $f$-DD, a similar generalization bound also applies to the Jeffereys divergence (or symmetrized KL divergence) [43] counterpart, which is simply the sum of KL divergence and reverse KL divergence (i.e. $\mathrm{D}_{\rm KL}(\hat{\mu}||\hat{\nu})+\mathrm{D}_{\rm KL}(\hat{\nu}||\hat{\mu})$). Furthermore, considering the fact that $\mathrm{D}_{\rm KL}(\nu||\mu)\leq\log(1+\chi^{2}(\nu||\mu))\leq\chi^{2}(\nu||\mu)$ [30], one might anticipate a similar bound for $\chi^{2}$-DD, which we defer to Appendix D.11.
This generalization bound suggests that when $r+r_{1}$ is small, not only are the first four terms (including the local Rademacher complexities) reduced, but it also causes $\mathcal{O}\left(\frac{1}{n}+\frac{1}{m}\right)$ to dominate the convergence rate of the bound. In practice, when empirical risk of the source domain is always minimized to zero (i.e. the realizable case), then $r_{1}$ itself may have a fast vanishing rate (e.g., $\mathcal{O}(1/n)$). In Appendix D.12, we provide a concrete example to further illustrate the superiority of localized $f$-DD.
6 Algorithms and Experimental Results
6.1 Domain Adversarial Learning Algorithm
In a practical algorithm, the hypothesis space consists of two components: the representation part, denoted as $\mathcal{H}_{\rm rep}=\{h_{\rm rep}:\mathcal{X}\to\mathcal{Z}\}$, where $\mathcal{Z}$ is the representation space (e.g., the hidden output of a neural network), and the classification part, denoted as $\mathcal{H}_{\rm cls}=\{h_{\rm cls}:\mathcal{Z}\to\mathcal{Y}\}$. Therefore, the entire hypothesis space is given by =$\mathcal{H}=\{h_{\rm cls}\circ h_{\rm rep}|h_{\rm rep}\in\mathcal{H}_{\rm rep},h_{\rm cls}\in\mathcal{H}_{\rm cls}\}$. The training objective in $f$-DAL [1] is